Let a total of 48 MHz of bandwidth be allocated to a particular FDD cellular telephone system, which uses two 36 kHz simplex channels to provide full duplex voice and control channels. Compute the number of channels available per cell if the system uses 8-cell reuse. If 1.44 MHz of the allocated spectrum is dedicated to control channels, determine an equitable distribution of control channels and voice channels in each cell.
Answer #1

Given total bandwidth = 48MHz = 48000KHz

Channel bandwidth

= 36KHz * 2 simplex channels

= 72KHz per duplex channels

Total no of channels =total bandwidth/channel bandwidth = 48000/72= 666.66= 667

Total no of channels =667

Number of cell reuse N=8

Total no of channels per cell = 667 /8 = 84

Spectrum allocated to control channels is 1.44MHz = 1440KHz

Total no of control channels = 1440/ 72 = 20 channels

Total no of voice channels available are 667-20= 647

For 8 cell resuse ,

No of control channels per cell = 20/8 = 3

No of voice channels per cell = 647/8 = 81

Therfore inequitable distribution of control channels is 3 per cell and voice channels is 81 per cell.